OpenStudy (anonymous):

If g(x) = 3(x − 2), find the value of x if g(x) = 6. 12 2.6 8 4 i got 4 but I'm not sure

OpenStudy (greatlife44):

actually set that equal to 6 and then solve for x

OpenStudy (greatlife44):

6 = 3(x-2) then distribute the 3 6 = 3x-6 add 6 to both sides 12 = 3x divide by 3 x = 4

OpenStudy (anonymous):

thank you so much :)

OpenStudy (greatlife44):

@hopemallard remember g(x) = 6 means like what y value the function means, it's the opposite of say g(6) which means that 6 is the x value

OpenStudy (greatlife44):

you had it down already lol =]